OpenStudy (anonymous):

Evaluate the expression. 9P0

OpenStudy (astrophysics):

\[_9P_0 = \frac{ 9! }{ 0! \times 9! }\] do you mean this?

OpenStudy (anonymous):

Yes, thank you. :)

OpenStudy (anonymous):

so my answer was 9 is that correct?

OpenStudy (astrophysics):

Nope, because 0! = 1 and 9!/9! cancel each other out hence your answer will be left as 1/(1x1) = 1

OpenStudy (anonymous):

1, thank you...it makes total sense now.

OpenStudy (astrophysics):

Np ^.^